in a recent election, the winning candidate had 2,700 more votes than the loser. if the total number of votes was 13,300, how many votes did the winner receive ?
Answer by ikleyn(52776)

.

One way to solve

    winning candidate + other candidate = 13300

        x             +   (x - 2700)    = 13300

       2x                               = 13300 + 2700 = 16000

        x                                              = 16000/2 = 8000


    ANSWER.  Winners received 8000 votes.


ANOTHER way to solve

You write the system of two equations


    X + Y = 13300

    X - Y =  2700


and add equations.


You will get the same answer for the winner.

Solved, explained and completed.


--------------

If you want to see a million similar solved problems, look into the lesson
    - Word problems that lead to a simple system of two equations in two unknowns
in this site.


Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Systems of two linear equations in two unknowns".

Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.